Assume you have the following data: H0: = 200, S = 30, n = 64, and X-bar= 218. Conduct a two-tailed hypothesis test at the .05 significance level.
Answer by stanbon(75887)

Assume you have the following data:
Ho: u = 200
Ha: u is not equal to 200
--------------------------

S = 30, n = 64, and X-bar= 218. Conduct a two-tailed hypothesis test at the .05 significance level.
t(218) = (218-200)/[30/sqrt(64)] = 18/[30/8] = 9/(15/8) = 72/15 = 4.8
---
p-value = 2*P(t > 4.8 when df = 63) = 2*tcdf(4.8,100,63) = 0.00001014
-=====
Conclusion: Since the p-value is less than 5%, reject Ho.
The test supports the belief that u is not equal to 200
at the 5% significance level.
=============
Cheers,
Stan H.
===============